OpenStudy (anonymous):

A thin rod of length L cm and a small particle lie on a line separated by a distance of A cm. If K is a positive constant and F is measured in Newtons the gravitational force btwn them is: F=K/((a(a+L))

OpenStudy (anonymous):

If a is increasing at the rate 2cm/min when a =15 and L=5, How fast is F decreasing?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

sorry i don't know a single thing about physics

OpenStudy (anonymous):

it isn't physics this is calc

OpenStudy (anonymous):

it is related rates

OpenStudy (anonymous):

oooooooooooh ok let me try to read carefully

OpenStudy (anonymous):

is this \[F=\frac{K}{a(a+L)}\]?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok so we want F' and K and L are fixed. so we take the derivative wrt time right?

OpenStudy (anonymous):

well we fisrt take the derivative assuming that l is fixed and a isn't

OpenStudy (anonymous):

since we see that a is changing

OpenStudy (anonymous):

no a is not fixed. K is a constant and L is a constant. a is changing and so is F

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

I got an answer but a real ugly one

OpenStudy (anonymous):

so i get \[F'=\frac{-K(2aa'+a'L)}{(a^2+aL)^2}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i see hwere i went wrong already

OpenStudy (anonymous):

now you know all the numbers so it should just be a substitution. i did not really use the "quotient rule' i used the fact that \[(\frac{1}{f})'=\frac{-f'}{f^2}\] and the chain rule

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

also i see that your answer will have a K in it right? because we don't know K we know L = 5, a i= 15, a' = 2

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

I still got the same ugly number

OpenStudy (anonymous):

i would use a calculator

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

let me try it and see

OpenStudy (anonymous):

i think i got \[-\frac{7K}{9000}\]

OpenStudy (anonymous):

I got -7k/900

OpenStudy (anonymous):

oh so maybe this is it

OpenStudy (anonymous):

I am always doubting myself

OpenStudy (anonymous):

i mean i am sure i got it. maybe i put in the numbers wrong, but this is what i computed \[-\frac{K(2\times 15\times 2+2\times 5)}{(15^2+15\times 5)^2}\]

OpenStudy (anonymous):

ya ok

OpenStudy (anonymous):

check to see that i put the right numbers in the right spot. i could have messed up

OpenStudy (anonymous):

so that must be the answer

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

k there is another part to the question

OpenStudy (anonymous):

I just need you to tell me if something i scorrect

OpenStudy (anonymous):

a simple one

OpenStudy (anonymous):

the next part says that L is decreasing by 2 cm so I guess a is constant and now L is varying?

hero (hero):

Related Rates suck. She's posting like all of the hardest questions. That's the last question in the section I bet

OpenStudy (anonymous):

Nope it isn't

OpenStudy (anonymous):

It is one of the first

OpenStudy (anonymous):

Sattelite is gone with the wind

hero (hero):

If that's the first question, I'd hate to see the last one

OpenStudy (anonymous):

It was a stupid question and he just ran off on me

OpenStudy (anonymous):

Ya I am stuck on one

OpenStudy (anonymous):

I have 4 more left to go

hero (hero):

You should go to bed. It's getting late

OpenStudy (anonymous):

ok now what? a is fixed and L is increasing??

OpenStudy (anonymous):

Oh ur back?????

OpenStudy (anonymous):

What happened to u?

OpenStudy (anonymous):

Well neways I used my own brain but I have another question and i will post it in another post

OpenStudy (anonymous):

how can L change? it is a length?

OpenStudy (anonymous):

if l is decreasing at the rate 2 cm /min when a =15 and and L=5 how fast is a increasing

OpenStudy (anonymous):

so I guess L is decreasing

OpenStudy (anonymous):

ok so now we have to start over and take the derivative wrt L right?

OpenStudy (anonymous):

I already solved this one. ya that is what i did

OpenStudy (anonymous):

u don't have to do it but i will post another question but can u pls come and help?

OpenStudy (anonymous):

Cherry please ;-)

OpenStudy (anonymous):

oh ok. i was hoping that a and L were not both moving, but i guess we could deal with that too if we had to. sure post and i will be there

OpenStudy (anonymous):

hey wait a sec

OpenStudy (anonymous):

ok i will wait

OpenStudy (anonymous):

A isn't moving in the second part?

OpenStudy (anonymous):

I don't think a is varying here. Do u think so?

OpenStudy (anonymous):

I solved it as a being a constant

OpenStudy (anonymous):

i don't know i can't see the question. we took the derivative the first time as if "a" was a variable and K and L were constants. if L is a function then we have to take the derivative differently

OpenStudy (anonymous):

A thin rod of length L cm and a small particle lie on a line separated by a distance of A cm. If K is a positive constant and F is measured in Newtons the gravitational force btwn them is: F=K/((a(a+L))

OpenStudy (anonymous):

but the way the first question was phrased it looked like L was constant

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

A) If a is increasing at the rate 2cm/min when a =15 and L=5, How fast is F decreasing?

OpenStudy (anonymous):

ok we did that right i believe

OpenStudy (anonymous):

B) if l is decreasing at the rate 2 cm /min when a =15 and and L=5 how fast is a increasing

OpenStudy (anonymous):

yes we did part a and this is part b

OpenStudy (anonymous):

i am not sure how a thin rod can be decreasing in length. does this make any sense?

OpenStudy (anonymous):

B) if L is decreasing at the rate 2 cm /min when a =15 and and L=5 how fast is a increasing

OpenStudy (anonymous):

it must be

OpenStudy (anonymous):

in this case you would have \[F'=\frac{-akL'}{(a^2+aL)^2}\] i believe

OpenStudy (anonymous):

ya ok so that is what i got. I made a a constant

OpenStudy (anonymous):

now \[L'=-2,a=15,L=5\] and replace

OpenStudy (anonymous):

whtvr it is fine. If i get it wrong then I will learn what to do for next time

OpenStudy (anonymous):

looks good to me. i think the idea is to introduce you to partial derivatives, where you hold one thing constant and take the derivative wrt the other variable. i think this looks right. post the next one and i will look.

OpenStudy (anonymous):

ok thanks u r the best!!!!